OpenStudy (anonymous):

Two of the top grossing concert tours were by a jazz band and a rock band. Together the two tours visited 183 cities. The jazz band had visited 97 cities more than the rock band. How many cities did each group visit. I cant really translate word problems too well can anyone show me how I would set this up and solve it?

OpenStudy (anonymous):

Let x be the no. of tours visited by rock band and then the jazz band's tours will be x+97, \[x+(x+97) =183\]You can solve the rest right

OpenStudy (katrinakaif):

Let j = jazz band Let r = rock band Both visited 183 cities in all j + r = 183 Jazz band visited 97 more than r = j = r+ 97 Substitute in the equation before giving r +97) + r = 183

OpenStudy (anonymous):

so would i subtract 97 from both sides?

OpenStudy (anonymous):

yes after that divide both sides by two as you would get 2x=86

OpenStudy (anonymous):

so x=43?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so now that i have 43 i substitute that for y? in the equation

OpenStudy (anonymous):

we assumed x to be the no. of tours rock band visited so the jazz band will have visited x+97=143 tours

OpenStudy (anonymous):

so i plug in 43+97=143?

OpenStudy (anonymous):

yes